OpenStudy (anonymous):

PLEASE HELP!!! Use the substitution method to solve the following system of equations. x + 2y - z = 2 3x + y - 2z = -15 3x + 4y - 3z = -2

OpenStudy (anonymous):

a) (7, -4, -1) b) (-7, -4, 1) c) (7, -4, 1) d) (-7, 4, -1)

OpenStudy (bahrom7893):

x = 2-2y+z substitute into the 2nd equation: 3(2-2y+z) + y - 3z = -15 solve for y in this one, and then substitute both into the last equation.

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

my final answer is b) (-7, -4, 1)

OpenStudy (bahrom7893):

Guessing won't work here... Let's try out B: -7+2(-4) -1 = -7 - 8 - 1 = -16 (the first equation is already wrong)

OpenStudy (anonymous):

no i didnt guess i actually worked it out

OpenStudy (bahrom7893):

There's a mistake in your solution somewhere then

OpenStudy (anonymous):

i guess i did it wrong tho

OpenStudy (anonymous):

ya i dont know this is hard

OpenStudy (bahrom7893):

Try d

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i tried and i think i messed up but imma just go with d

OpenStudy (anonymous):

i tried all the equations

OpenStudy (bahrom7893):

Verify it using my method

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

IS THE ANSWER D????

OpenStudy (anonymous):

or could you check for me? and show me how because im confused

OpenStudy (anonymous):

ive been stuck on this question for over 2 hours and i need to turn this assignment in asap could you help?

zepdrix (zepdrix):

\[\large 3x+y-2z=-15\]Rewriting the second equation, moving some stuff around, gives us, \[\large \color{green}{y=-15-3x+2z}\] \[\large x+2\color{green}{y}-z=2\]Plugging this into the first equation gives us,\[\large x+2\color{green}{\color{green}{(-15-3x+2z)}}-z=2\] Now we've successfully written one of our equations in terms of only 2 variables. You're suppose to solve this using substitution like this? Wow this is going to be a long process :D lol

zepdrix (zepdrix):

Yah plugging in the options is probably a better idea.. hmm

OpenStudy (anonymous):

i think the answer is d (-7,-4,1)

OpenStudy (anonymous):

i worked it out and thats what i got.

zepdrix (zepdrix):

d says (-7, 4, -1), not (-7, -4, -1). You're just looking for someone to check your work I guess..? :o Yah good job. That's the correct answer.

OpenStudy (anonymous):

my bad i accidentally put the negative on the 4

OpenStudy (anonymous):

@zepdrix so the correct answer is d (-7,4,1)

zepdrix (zepdrix):

Yes, the correct answer is d, you keep writing it incorrectly though lol :) (-7, 4, -1) not (-7, 4, 1)

OpenStudy (anonymous):

sorry haha its late. thank you though

zepdrix (zepdrix):

XD